log3(x) + log3(6x)=3
Answer by stanbon(75887)

log3(x) + log3(6x)=3
-------
log3[x*6x) = 3
---
log3(6x^2) = 3
----
6x^2 = 3^3
6x^2 = 27
x^2 = 27/6
-----
x = [3sqrt(3)/6]
-----
x = sqrt(3)/2
=========================
Cheers,
Stan H.
=========================